OpenStudy (anonymous):

Pablo rolled a standard die twenty-four times. He got a 1 three times. How does this result compare to the expected results? Explain.

OpenStudy (anonymous):

probability of rolling a one is \(\frac{1}{6}\) so you expect \(24\times \frac{1}{6}=4\) ones in 24 rolls

Parth (parthkohli):

Well, theoretically, we should have gotten a three in six moves. \(\Large \color{MidnightBlue}{\Rightarrow 1:6=x:24 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = 4 }\) We were expecting a three 4 times theoretically.

OpenStudy (anonymous):

you expect 1/6 of all rolls to be a 1.

Directrix (directrix):

From @Will He rolled it 24 times, and there are six possible results of a roll. So dividing 24 by 6 means that each number should have come up four times. His result was one fewer than what was expected.